OpenStudy (anonymous):

can someone help me with this cotxsec^4(x)=cotx+2tanx+tan^3(x)

OpenStudy (science0229):

Let's start with the left side.

OpenStudy (science0229):

can you express the left side in terms of sin and cos?

hartnn (hartnn):

alternatively, if you want to do it without converting everything to sin and cos, you can just start by writing \(\sec^4 x = (\sec^2 x)^2\) using the pythagorean identity for sec^2 and then expanding :) this method could be faster :)

OpenStudy (science0229):

I'll do both methods, even though I have to agree with @hartnn (It's been a while since I played with these identity problems)

OpenStudy (science0229):

The identity is \[1+\tan^2=\sec^2\]

OpenStudy (science0229):

I'll skip writing x for my own convenience

OpenStudy (anonymous):

hold on, ive been working on it let me show you what i got

OpenStudy (science0229):

Sure.

OpenStudy (anonymous):

1+tan^3(x)=cos^2(x)+2sinx/2cosx+tan^3(x)

OpenStudy (science0229):

How did you get to there? you can just say it simply

OpenStudy (anonymous):

how

OpenStudy (science0229):

\[\cot(\sec^2)^2=\cot(1+\tan^2)^2=\cot(1+2\tan^2+\tan^4)=\cot+2\tan+\tan^3\]

OpenStudy (science0229):

That's the left side

OpenStudy (science0229):

which is equal to the right side the proof is complete

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im confused

hartnn (hartnn):

in that proof, we used the fact that since, tan x = 1/cot x cot x * tan x =1

OpenStudy (science0229):

oh sorry forgot to explain that

hartnn (hartnn):

4 steps are shown as the proof...which step are you confused at ?

OpenStudy (anonymous):

are you guys going based on the original problem or the simplified one i put?

OpenStudy (science0229):

the original problem

OpenStudy (science0229):

notice that \[\cot*\sec^4=\cot(\sec^2)^2\]

OpenStudy (anonymous):

but no where in the original problem does it say cotx*tanx it says cotx+2tanx

OpenStudy (science0229):

from third to fourth step...\[\cot(1+2\tan^2+\tan^4)=\cot*1+\cot*2\tan^2+\cot*\tan^4=\cot+2\tan*\tan*\cot+\]\[\cot*\tan*\tan^3=\cot+2\tan+\tan^3\]

OpenStudy (science0229):

do you have any questions?

OpenStudy (anonymous):

nope let me just go over it right quick

OpenStudy (science0229):

sure. take as much time as needed

OpenStudy (anonymous):

im so sorry but im just not getting it

OpenStudy (science0229):

I'm really sorry, but I have to go.

OpenStudy (anonymous):

no problem thanks though